OpenStudy (anonymous):

help!? pleaseeee :) Simplify (6x2+ 11x – 3) + (2x2– 17x – 4) 8x2+ 6x – 1 8x2– 6x – 7 8x2+ 6x + 1 8x2– 6x + 7 How do you simplify these equations?

OpenStudy (anonymous):

add like terms

OpenStudy (raden_zaikaria):

6x2+ 11x – 3 + 2x2– 17x – 4

OpenStudy (anonymous):

(6x2+ 11x – 3) + (2x2 – 17x – 4) 8x^2 -4x -7

OpenStudy (anonymous):

\[6x^2+ 11x – 3 + 2x^2– 17x – 4\]

OpenStudy (anonymous):

B I think

OpenStudy (anonymous):

Yeah Its B

OpenStudy (anonymous):

Setting it up like this makes it alot easier: (6x2+ 11x – 3) (2x2– 17x – 4)

OpenStudy (anonymous):

6x2 + 2x2 = 8x2 11x - 17x = -6x -3 - 4 = -7

OpenStudy (raden_zaikaria):

6x2+ 2x2=8x2 and 11x – 17x = -6x and – 3 – 4 = -7

OpenStudy (anonymous):

(6x2+ 11x – 3) + (2x2 – 17x – 4) 8x^2 -6x -7 sorry

OpenStudy (raden_zaikaria):

that ansmwer is 8x2-6x-7

OpenStudy (anonymous):

Thank you! everyone! :) I get it now :)

OpenStudy (anonymous):

Any time

OpenStudy (raden_zaikaria):

welcome from malaysia hehe

OpenStudy (anonymous):

Simplify 2x2(4x2+ 9x – 13) 6x4+ 18x3– 26x2 8x4+ 18x3– 26x2 8x2+ 18x – 26 6x2+ 18x – 26

OpenStudy (anonymous):

would the answer be D? or A...

OpenStudy (anonymous):

none

OpenStudy (anonymous):

oh........ why?

OpenStudy (anonymous):

\[2x^2(4x^2+ 9x – 13)\] now we are multiplying

OpenStudy (anonymous):

so \[2x^2\] should multiply each item in the bracket

OpenStudy (anonymous):

hmm, so C?

OpenStudy (anonymous):

almost there but notice that even the power changes \[2x^2(4x^2)=8x^{4}\]

OpenStudy (anonymous):

the powers add each other\[x^a(x^b)=x^{a+b}\]

OpenStudy (anonymous):

oh. ok. I wasnt sure if i was supposed to change the powers as well.

OpenStudy (anonymous):

Which of the following represents the area of a rectangle whose length is x – 7 and whose width is x + 12? x2+ 5x – 84 x2– 84 x2– 5x – 84 x2– 84x + 5 IS C THE CORRECT ANSWER?

OpenStudy (anonymous):

lol...

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

so did you get the answer to the previous one

OpenStudy (anonymous):

Yes I did.

OpenStudy (anonymous):

you shud ask a new question as a new question

OpenStudy (anonymous):

I'm not 100% sure on this question :l

OpenStudy (anonymous):

ok,

OpenStudy (anonymous):

I'll close this question and post a new one so i can give u a medal.